OpenStudy (anonymous):

Two angles are supplementary. The larger angle is 20 degrees less than 7 times the smaller angle. Find the measures of the angles.

OpenStudy (anonymous):

Let x represent the smaller angle Let y represent the larger angle y = 7x - 20 180 = x + y

OpenStudy (anonymous):

you can use substitution or elimination to solve it

OpenStudy (anonymous):

thankyuu :D

OpenStudy (anonymous):

in this case i would use substitution since its already done for us y = 7x - 20 (1) 180 = x + y (2) substitute (1) into (2) 180 = x + y 180 = x + (7x - 20) 180 = x + 7x - 20 180 + 20 = 8x *Move 20 over to the left side* 200/8 = 8x/ 8 25 = x *The smaller angle is 25 degrees substitute x = 25 degrees into any equation :) (i chose equation (1) ) 180 = x + y 180 = 25 + y 180 - 25 = y 155 = y * The larger angle is 155 degrees Write your therefore statement :D